# Weylian.bib

### v2

@@ 2.2 Abundant superintegrable systems
We now introduce the concept of abundant superintegrable systems. For the following definition, we will denote the Ricci tensor of the Riemannian metric $g$ by $\mathrm{Ric}$,
@@
We may omit the subscript $g$, if the underlying metric is clear.
The following definition is based on the structural equations established in [KSV2023, KSV2024, KSV2024_bauhaus]. The terminology follows [CV2025].
@@
Let $(M,g)$ be a conformally flat Riemannian manifold of dimension $n\geq 2$, endowed with a totally symmetric and trace-free tensor field $S\in\Gamma(\mathrm{Sym}^{3}_{\circ}(T^{*}M))$ and a function $t\in\mathcal{C}^{\infty}(M)$ such that [^1] Compared to [KSV2024_bauhaus], two typographical errors have been corrected in (5fb) and (5fd), involving the replacement of two erroneous signs.
@@
    $\nabla_{i}\bar{t}_{j}$ $=3\,\mathsf{P}_{ij}+\frac{1}{3(n-2)}\left({S_{i}}^{ab}S_{jab}+\frac{n-6}{2(n-1)(n+2)}\,|S|_{g}^{2}\,g\right)$
    $\qquad\qquad\qquad+\frac{1}{3}\left(\nabla_{i}t\nabla_{j}t-\frac{1}{2}|\mathrm{grad}_{g}t|_{g}^{2}\,g\right)$ (5a)
@@
    00 0 $=\Pi_{\circ}\left(g^{ab}S_{aik}S_{bjl}-g^{ab}S_{ail}S_{bjk}\right).$ (5c)
@@
    $\nabla_{l}S_{ijk}$ $=\Pi_{(ijk)\circ}\left[-\frac{2}{3}S_{ijk}t_{l}+2t_{i}S_{jkl}+\Xi_{ij}g_{kl}\right]$ (5fa)
    $\nabla_{k}\Xi_{ij}$ $=\frac{1}{3}S_{ijk}\,\,|S|_{g}^{2}+\frac{4}{3}\Pi_{(ijk)\circ}\Xi_{jk}t_{i}+\frac{4}{3}\Pi_{(ij)\circ}g_{jk}\beta_{i}$ (5fb)
@@
    $\hskip 170.71652pt+\frac{1}{2}\nabla_{k}R-Rt_{k}$ (5fc)
@@
    $\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad+\frac{1}{2}g_{ij}\left[\frac{1}{3}\,|S|_{g}^{2}+\frac{3}{2}R\right]$ (5fd)
    where we introduce the $1$-form $\beta\in\Omega^{1}(M)$, defined by
    $\beta_{k}$ $=S_{kab}\Xi^{ab}\,.$
@@
    The tensor fields $\Xi,\tau\in\Gamma(\mathrm{Sym}^{2}_{\circ}(T^{*}M))$ in Definition 2 will in the following be considered as mostly technical, auxiliary objects. However, we mention that $\tau$ has a deeper significance for distinguishing properly and conformally superintegrable systems, cf. [KSV2024, BLMS24] and also [GKL2024].
@@ 3 Abundant structures and Weylian manifolds
    Having introduced abundant structures and Weylian manifolds, we are now going to relate these two structures. Recall that for an exact Weyl structure $(M,c,D)$, the $1$-form $\theta_{g}$ in (1) is exact for any $g\in c$. We thus have $\theta_{g}=d\Theta_{g}$ for some function $\Theta_{g}\in\mathcal{C}^{\infty}(M)$. Analogously, consider the corresponding Weylian structure $(M,\Phi)$. For $(g,\varphi)\in\Phi$, the $1$-form $\varphi$ is also exact, and $\varphi=2d\phi$ for some function $\phi$, c.f. Remark 1. With this terminology, we state the following lemma, which allows us to define abundant structures on Weylian manifolds. Its proof is implied by [KSV2024].
@@
    Let $(M,\Phi)$ be an exact Weylian structure. Let $(g,\varphi)\in\Phi$ and assume that $(M,g)$ admits an abundant structure encoded in the tensor field $S\in\Gamma(\mathrm{Sym}^{3}_{\circ}(T^{*}M))$ and the function $t\in\mathcal{C}^{\infty}(M)$. Next, let $(g^{\prime},\varphi^{\prime})$ be another representative of the Weylian metric $\Phi$.
    Then $(M,g^{\prime})$ carries the abundant structure encoded in the structure tensor $S^{\prime}=g^{\prime}g^{-1}S$ and the function $t^{\prime}=t-3(\varphi-\varphi^{\prime})$.
    Note that $g^{\prime}g^{-1}S$ here means that we first raise, then lower one index using the metrics $g$ and $g^{\prime}$, respectively.
    Proof.
@@
    We observe that $\hat{S}=g^{-1}S\in\Gamma(\mathrm{Sym}^{2}_{\circ}(T^{*}M)\otimes TM)$ and $\hat{S}=(g^{\prime})^{-1}S^{\prime}\in\Gamma(\mathrm{Sym}^{2}_{\circ}(T^{*}M)\otimes TM)$ coincide. We also note that in the lemma, $\frac{1}{3}t$ and $\varphi$ transform into $\frac{1}{3}t^{\prime}$ and $\varphi^{\prime}$, respectively, in an analogous manner. The equation $t-3\varphi=t^{\prime}-3\varphi^{\prime}$ hence yields a conformal invariant. We are now ready to define abundant structures on Weylian manifolds.
@@
    We call a triple $(M,\Phi,\hat{S})$ consisting of a smooth manifold $M$ with Weylian metric $\Phi$ as well as a (2,1)-tensor field $\hat{S}\in\Gamma(\mathrm{Sym}^{3}_{\circ}(T^{*}M)\otimes TM)$ a natural abundant Weylian structure if $S=g\hat{S}\in\Gamma(\mathrm{Sym}^{3}_{\circ}(T^{*}M))$ is trace-free and satisfies the conditions in Definition 2 with $t=3\varphi$.
    We introduce natural abundant Weyl structures and natural abundant Weylian manifolds accordingly, in the obvious way. Note that in the definition, we suppress for the sake of conciseness the irrelevant constant that may be added to $t$ or $\varphi$ (i.e. we identify structures that only differ by such a constant).
@@
    Compare Definition 4 to the definition of c-superintegrable systems, c.f. Definition 3.9 in [KSV2024]. We note that our focus here is on the case of abundant structures, but that the underpinning ideas are nontheless of a general nature. Indeed, we observe the fundamental difference that, in the reference, a c-superintegrable system is defined on a merely conformal structure $(M,c)$, i.e. a smooth manifold with a conformal class of metrics. The discussion in the current section has revealed the underpinning Weylian geometry. In [KSV2023], this underlying Weylian structure exists implicitly, c.f. Section 6 of [KSV2024]. For $n\geq 3$, it is indirectly inferred, by Eq. (6.6) of [KSV2023], from a conformal class of abundant superintegrable systems.
    We may reinterpret this as having subsumed the invariant difference function $\mathfrak{t}=t-3\varphi$ bridging between the representative $(g,\varphi)\in\Phi$ of the Weylian metric and the abundant structure $(M,g,S,t)$. Indeed, if $\mathfrak{t}\neq 0$, we may replace $\Phi$ be a new Weylian metric $\Phi^{\prime}=[(g,\varphi+\frac{1}{3}\mathfrak{t})]$. In terms of the corresponding Weyl structure this presents itself as a conformal change of the Weylian connection. This observation now allows us to extend abundant structures to pre-fixed Weylian manifolds.
@@
    Let $(M,\Phi)$ be a Weylian structure. We call $(\hat{S},\mathfrak{t})$ consisting of a (2,1)-tensor field $\hat{S}\in\Gamma(\mathrm{Sym}^{2}_{\circ}(T^{*}M)\otimes TM)$ and a function $\mathfrak{t}\in\mathcal{C}^{\infty}(M)$ an abundant structure on $(M,\Phi)$ if $(M,[(g,\varphi+\frac{1}{3}\mathfrak{t})],\hat{S})$ is an abundant Weylian manifold.
@@
      If $(\hat{S},\mathfrak{t})$ is an abundant structure on the Weylian manifold $(M,\Phi)$, then $(M,[(g,\varphi+\mathfrak{t})],\hat{S})$ is a natural abundant Weylian manifold, $(g,\varphi)\in\Phi$.
@@
    For part (i), observe that the structural equations (5) and (5f), respectively, involve only derivatives of the function $t$.
@@ 3.1 Structural equations for $n\geq 3$
    Our aim is now to give a “representative-free” characterisation of abundant structures on Weylian manifolds. To this end, we need to establish some terminology for a given Weyl structure $(M,g,D)$. In order to do this efficiently, we introduce some terminology. We denote by $\Pi_{\mathrm{Sym}^{3}_{0}}$ the projection onto the trace-free totally symmetric component, and we introduce conformally adapted operators $\mathscr{H}:\mathcal{C}^{\infty}(M)\to\Gamma(\mathrm{Sym}^{2}_{\circ}(T^{*}M))$,
@@
    $\mathcal{P}$ $=-\frac{1}{9(n-2)}\,\mathring{\mathscr{S}}$ (5g)
    $\mathcal{L}$ $=-\frac{2}{9}\frac{3n+2}{n+2}\,|S|^{2}$ (5h)
    $D_{X}\hat{S}$ $=\mathfrak{A}(X,\cdot,\cdot)$ (5i)
    $\mathrm{Weyl}^{\hat{S}}$ $=0$ (5j)
    Note that (5g)–(5j) are conformally invariant conditions, i.e. that they hold true also after a conformal rescaling.
@@
    The conditions (5g)–(5j) are obviously equivalent to (5). Hence, for each $(g,\phi)\in\Phi$, $(M,g,g\hat{S},t)$, $t=\frac{1}{3}\phi$, defines an abundant manifold. Hence, $(M,\Phi,\hat{S})$ defines an abundant Weylian manifold. ∎
@@ 3.3 Semi-Weyl manifolds
    Re-examining the geometric structure of abundant superintegrable systems, we have explicitly described the underlying Weyl structure. This allowed us to encode conformal classes of abundant structures in a totally symmetric and trace-free tensor field $S\in\Gamma(\mathrm{Sym}^{3}_{\circ}(T^{*}M))$ on a Weyl structure $(M,c,D)$. Consider an abundant structure $(M,g,S,t)$. Then the corresponding natural Weyl structure is given by the representative $(g_{0},0)$ where $g_{0}=e^{-\frac{2}{3}t}$. Indeed, recalling Lemma 1, a representative of the Weylian metric is given by
@@
    where $\Omega$ is a smooth function. Asking that $dt-d\ln|\Omega|=0$ (we recall that the Weylian manifold is exact), we obtain $\Omega^{2}=\exp(-\frac{2}{3}t)$ (modulo multiplication by an irrelevant constant). It follows that, letting $g_{0}:=e^{-\frac{2}{3}t}g$,
@@
    We remark that, in the terminology of [KSV2024], the metric $g_{0}$ is the metric of the “system in standard scale” associated to the system corresponding to the abundant structure $(M,g,S,t)$ of the theorem.
    Changing our vantage point, we will now explain how this allows us to conceptualise the abundant structure as a semi-Weyl manifold. We find the following definition in [Matsuzoe2010].
@@
    An abundant structure $(M,g,S,t)$ has an underlying semi-Weyl structure. In particular, denoting by $\nabla^{0}$ the Levi-Civita connection of $\exp(-\frac{2}{3}t)g$, let
@@
    We have hence found that the data of $(M,g,S,t)$ can alternatively be encoded in the semi-Weyl structure $(M,g,\bar{D},-\frac{2}{3}dt)$ where we identify abundant structures $(g,S,t)\sim(g,S,t^{\prime})$ when $t-t^{\prime}$ is a constant.
    We also mention that this endows $(M,g_{0},g_{0}g^{-1}S)$ with the structure of a statistical manifold, cf. [Matsuzoe2010].

### v3

@@ 2.2.1 Abundant structures
We proceed to state the definition of an abundant structure, for which we introduce the following notation. We denote the Ricci tensor of the Riemannian metric $g$ by $\mathrm{Ric}$,
@@
We may omit the subscript $g$, if the underlying metric is clear. Also, for better legibility, we abbreviate $t_{i}:=\nabla_{i}t$ in the following, leading to the natural equality $\bar{t}_{i}=\nabla_{i}t=t_{i}$.
In the remainder of the paper, we furthermore adopt Einstein’s summation convention for repeated pairs of (upper and lower) indices.
@@
Let $(M,g)$ be a conformally flat Riemannian manifold of dimension $n\geq 2$, endowed with a totally symmetric and trace-free tensor field $S\in\Gamma(\mathrm{Sym}^{3}_{\circ}(T^{*}M))$ and a function $t\in\mathcal{C}^{\infty}(M)$ such that [^3] Compared to [KSV2024_bauhaus], two typographical errors have been corrected in (9b) and (9d), involving the replacement of two erroneous signs.
@@
    $\nabla^{2}_{ij}t$ $=3\,\mathsf{P}_{ij}+\frac{1}{3(n-2)}\left({S_{i}}^{ab}S_{jab}+\frac{n-6}{2(n-1)(n+2)}\,|S|^{2}\,g\right)$
    $\qquad\qquad\qquad+\frac{1}{3}\left(t_{i}t_{j}-\frac{1}{2}|\mathrm{grad}_{g}t|^{2}\,g\right)$ (8a)
@@
    00 0 $=\Pi_{\circ}\left(g^{ab}S_{aik}S_{bjl}-g^{ab}S_{ail}S_{bjk}\right)$ (8c)
@@
    $\nabla_{l}S_{ijk}$ $=\Pi_{(ijk)\circ}\left[-\frac{2}{3}S_{ijk}t_{l}+2t_{i}S_{jkl}+\Xi_{ij}g_{kl}\right]$ (9a)
    $\nabla_{k}\Xi_{ij}$ $=\frac{1}{3}S_{ijk}\,\,|S|_{g}^{2}+\frac{4}{3}\Pi_{(ijk)\circ}\Xi_{jk}t_{i}+\frac{4}{3}\Pi_{(ij)\circ}g_{jk}\beta_{i}$ (9b)
@@
    $\hskip 170.71652pt+\frac{1}{2}\nabla_{k}R-Rt_{k}$ (9c)
@@
    $\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad+\frac{1}{2}g_{ij}\left[\frac{1}{3}\,|S|_{g}^{2}+\frac{3}{2}R\right]$ (9d)
    where for better legibility we introduce tensor fields $\Xi\in\Gamma(\mathrm{Sym}^{2}_{\circ}(T^{*}M))$, $\tau\in\Gamma(\mathrm{Sym}^{2}_{\circ}(T^{*}M))$ and the $1$-form $\beta\in\Omega^{1}(M)$, defined by
    $\Xi_{ij}$ $=\nabla^{a}S_{aij}+\frac{2}{3}\,S_{ija}t^{a}\,,$
    $\tau_{ij}$ $=\frac{2}{3}\,(\nabla^{2}_{ij}t-\tfrac{1}{n}\,\Delta t\,g_{ij})-\frac{2}{3}\,S_{ija}t^{a}+\frac{1}{3}\,\Xi_{ij}-\frac{8}{9}(t_{i}t_{j}-\tfrac{1}{n}\,t^{a}t_{a}\,g_{ij})\,,$
    $\beta_{k}$ $=S_{kab}\Xi^{ab}\,,$
    where $\Delta$ is the Laplace-Beltrami operator.
@@
Note that the equations in Definition 3 depend on $t$ only via $t_{i}=\bar{t}_{i}$. Next, observe that the tensor fields $\Xi,\tau\in\Gamma(\mathrm{Sym}^{2}_{\circ}(T^{*}M))$ are mainly introduced as auxiliary objects, and indeed we shall use them as mostly technical definitions. However, we comment that $\tau$ has a deeper significance. In fact it distinguishes properly and conformally superintegrable systems, cf. [KSV2024, BLMS24] and also [GKL2024]: if $\tau_{ij}=0$, then the abundant conformally superintegrable system admits proper integrals of the motion (rather than just on the zero locus of $H$). For completeness, we remark that for $n\geq 3$, this tensor is given by
[…]
@@
This and the next definition follow [CV2025], cf. [KSV2024, KSV2024_bauhaus].
Definition 5.
Two abundant structures $(g,S,t)$ and $(g^{\prime},S^{\prime},t^{\prime})$ on the same smooth manifold are said to be equivalent if $dt^{\prime}=dt$.
@@ 3 Abundant structures and Weylian manifolds
Having introduced abundant structures and Weylian manifolds, we are now going to relate these two structures. Recall that for an exact Weyl structure $(M,c,D)$, the $1$-form $\theta_{g}$ in (1) is exact for any $g\in c$. We thus have $\theta_{g}=d\Theta_{g}$ for some function $\Theta_{g}\in\mathcal{C}^{\infty}(M)$. Analogously, consider the corresponding Weylian structure $(M,\Phi)$. For $(g,\varphi)\in\Phi$, the $1$-form $\varphi$ is also exact, and $\varphi=d\phi$ for some function $\phi$, cf. Remark 1. We say that $\Theta_{g}$ and $\phi$ are potentials for $\theta_{g}$ and $\varphi$, respectively. With this terminology, we state the following lemma, which allows us to define abundant structures on Weylian manifolds. Its proof is implied by [KSV2024].
@@
Let $(M,\Phi)$ be an exact Weylian structure. Let $(g,\varphi)\in\Phi$ and $\phi$ a potential for $\varphi$, $\varphi=d\phi$. Assume that $(M,g)$ admits an abundant structure encoded in the tensor field $S\in\Gamma(\mathrm{Sym}^{3}_{\circ}(T^{*}M))$ and the function $t\in\mathcal{C}^{\infty}(M)$. Next, let $(g^{\prime},\varphi^{\prime})$ be another representative of the Weylian metric $\Phi$, with potential $\phi^{\prime}$, $\varphi^{\prime}=d\phi^{\prime}$.
Then $(M,g^{\prime})$ carries the abundant structure encoded in the structure tensor $S^{\prime}=g^{\prime}g^{-1}S$ and the function $t^{\prime}=t-3(\phi-\phi^{\prime})$.
Note that $g^{\prime}g^{-1}S$ here means that we first raise, then lower one index using the metrics $g$ and $g^{\prime}$, respectively. We also remark that $dt^{\prime}=dt-3(\varphi-\varphi^{\prime})$, and hence the abundant structure is determined by $\varphi^{\prime}-\varphi$ up to equivalence in the sense of Definition 5. Conversely, if $g^{\prime}=g$ and $t^{\prime}-t$ is constant, it must hold that $\varphi^{\prime}=\varphi$. The definite article for the abundant structure in the last sentence of the lemma is therefore justified.
Proof of Lemma 1.
@@
We observe that $\hat{S}=g^{-1}S\in\Gamma(\mathrm{Sym}^{2}_{\circ}(T^{*}M)\otimes TM)$ and $\hat{S}=(g^{\prime})^{-1}S^{\prime}\in\Gamma(\mathrm{Sym}^{2}_{\circ}(T^{*}M)\otimes TM)$ coincide. We also note that in the lemma, $t$ and $\varphi$ transform into $t^{\prime}$ and $\varphi^{\prime}$, respectively, in an analogous manner. The equation $\bar{t}-3\varphi=\bar{t}^{\prime}-3\varphi^{\prime}$ hence yields a conformal invariant, where $\bar{t}=dt,\bar{t}^{\prime}=dt^{\prime}$. It seems that this invariant was first observed in [Capel_phdthesis] in the case $n=3$. We are now ready to define abundant structures on Weylian manifolds.
@@
We call a triple $(M,\Phi,\hat{S})$ consisting of a smooth manifold $M$ with Weylian metric $\Phi$ as well as a (2,1)-tensor field $\hat{S}\in\Gamma(\mathrm{Sym}^{3}_{\circ}(T^{*}M)\otimes TM)$ a natural abundant Weylian structure if $S=g\hat{S}\in\Gamma(\mathrm{Sym}^{3}_{\circ}(T^{*}M))$ is trace-free and satisfies the conditions in Definition 3 with $dt=3\varphi$.
We introduce natural abundant Weyl structures and natural abundant Weylian manifolds accordingly, in the obvious way. Note that in the definition, we suppress for the sake of conciseness the irrelevant constant that may be added to $t$, cf. Definition 5.
@@
Compare Definition 6 to the definition of c-superintegrable systems, cf. Definition 3.9 in [KSV2024]. We note that our focus here is on the case of abundant structures, but that the underpinning ideas are nontheless of a general nature. Indeed, we observe the fundamental difference that, in the reference, a c-superintegrable system is defined on a merely conformal structure $(M,c)$, i.e. a smooth manifold with a conformal class of metrics. The discussion in the current section describes the underpinning Weylian geometry. In [KSV2023], this underlying Weylian structure exists implicitly, cf. Section 6 of [KSV2024]. For $n\geq 3$, it is indirectly inferred, by Eq. (6.6) of [KSV2023], from a conformal class of abundant superintegrable systems.
We may reinterpret this as having subsumed the invariant difference function $\mathfrak{t}=t-3\phi$ bridging between the representative $(g,\varphi)\in\Phi$ of the Weylian metric and the abundant structure $(M,g,S,t)$. Indeed, if $\mathfrak{t}\neq 0$, we may replace $\Phi$ by a new Weylian metric $\Phi^{\prime}=[(g,\varphi+\frac{1}{3}\,d\mathfrak{t})]$. In terms of the corresponding Weyl structure this presents itself as a conformal change of the Weylian connection. This observation now allows us to extend abundant structures to pre-fixed Weylian manifolds.
@@
Let $(M,\Phi)$ be a Weylian structure. We call $(\hat{S},\mathfrak{t})$ consisting of a (2,1)-tensor field $\hat{S}\in\Gamma(\mathrm{Sym}^{2}_{\circ}(T^{*}M)\otimes TM)$ and a function $\mathfrak{t}\in\mathcal{C}^{\infty}(M)$ an abundant structure on $(M,\Phi)$ if $(M,[(g,\varphi+\frac{1}{3}\,d\mathfrak{t})],\hat{S})$ is an abundant Weylian manifold.
@@
    If $(\hat{S},\mathfrak{t})$ is an abundant structure on the Weylian manifold $(M,\Phi)$, then $(M,[(g,\varphi+\frac{1}{3}\,d\mathfrak{t})],\hat{S})$ is a natural abundant Weylian manifold, $(g,\varphi)\in\Phi$.
@@
For part (i), observe that the structural equations (8) and (9), respectively, involve only derivatives of the function $t$. ∎
Recall from Remark 1 that a Weyl structure can always be modified by a change of the connection $D\to D+\alpha\otimes\mathrm{id}$ for a $1$-form $\alpha\in\Omega^{1}(M)$. We see the analogous freedom in the second part of Theorem 1: we can realign the Weyl structure relative to the superintegrable system as long as we preserve exactness of the Weyl structure. The first part of the theorem, however, tells us that, up to the equivalence stated in Definition 5, there is a unique natural base alignment of the Weyl structure relative to the superintegrable system, namely the one that makes $\mathfrak{t}$ a constant.
@@ 3.1 Structural equations for $n\geq 3$
Our aim is now to give a representative-free characterisation of abundant structures on Weylian manifolds. To this end, we need to establish some terminology for a given Weyl structure $(M,g,D)$. We denote by $\Pi_{\mathrm{Sym}^{3}_{0}}$ the projection onto the trace-free totally symmetric component, and we introduce conformally adapted operators $\mathscr{H}:\mathcal{C}^{\infty}(M)\to\Gamma(\mathrm{Sym}^{2}_{\circ}(T^{*}M))$,
@@
$\mathcal{P}$ $=-\frac{1}{9(n-2)}\,\mathring{\mathscr{S}}$ (11)
$\mathcal{L}$ $=-\frac{2}{9}\frac{3n+2}{n+2}\,|S|^{2}$ (12)
$D_{X}\hat{S}$ $=\mathfrak{A}(X,\cdot,\cdot)$ (13)
$\mathrm{Weyl}^{\hat{S}}$ $=0$ (14)
Note that (11)–(14) are conformally invariant conditions, i.e. that they hold true also after a conformal rescaling.
@@
The conditions (11)–(14) are obviously equivalent to (8). Hence, for each $(g,\varphi=d\phi)\in\Phi$, $(M,g,g\hat{S},t)$, $t=\frac{1}{3}\phi$, defines an abundant manifold. Hence, $(M,\Phi,\hat{S})$ defines an abundant Weylian manifold. ∎
@@ 3.3 Semi-Weyl manifolds
Re-examining the geometric structure of abundant superintegrable systems, we have explicitly described the underlying Weyl structure. This allowed us to encode, on a Weyl structure $(M,c,D)$, conformal classes of abundant structures in an invariant tensor field $\hat{S}\in\Gamma(\mathrm{Sym}^{2}_{\circ}(T^{*}M)\otimes TM)$, corresponding for $g\in c$ to a totally symmetric and trace-free tensor field $S:=g\hat{S}\in\Gamma(\mathrm{Sym}^{3}_{\circ}(T^{*}M))$. Consider an abundant structure $(M,g,S,t)$. Then the corresponding natural Weyl structure is given by the representative $(g_{0},0)$ where $g_{0}=e^{-\frac{2}{3}t}$. Indeed, recalling Lemma 1, a representative of the Weylian metric is given by
@@
where $\Omega$ is a smooth function. Asking that $dt-d\ln|\Omega|=0$ (we recall that the Weylian manifold is exact), we obtain $\Omega^{2}=\exp(-\frac{2}{3}t)$, modulo multiplication by an irrelevant constant. It follows that, letting $g_{0}:=e^{-\frac{2}{3}t}g$,
@@
We remark that, in the terminology of [KSV2024], the metric $g_{0}$ is the metric of the “system in standard scale” associated to the system corresponding to the abundant structure $(M,g,S,t)$ of the theorem. Changing our vantage point, we will now explain how this allows us to conceptualise the abundant structure as a semi-Weyl manifold. We find the following definition in [Matsuzoe2010].
@@
An abundant structure $(M,g,S,t)$ has an underlying semi-Weyl structure $(M,g,\bar{D},-\frac{2}{3}\,dt)$. In particular, denoting by $\nabla^{0}$ the Levi-Civita connection of $\exp(-\frac{2}{3}t)g=:g_{0}$, let
@@
We have hence found that the data of $(M,g,S,t)$ can alternatively be encoded in the semi-Weyl structure $(M,g,\bar{D},-\frac{2}{3}dt)$ where we identify abundant structures $(g,S,t)\sim(g,S,t^{\prime})$ when $t-t^{\prime}$ is a constant, cf. Definition 5. We also mention that this endows $(M,g_{0},g_{0}g^{-1}S)$ with the structure of a statistical manifold, cf. [Matsuzoe2010].
We conclude the paper with two examples.
Example 1.
Consider the Riemannian metric $g=\sum_{i=1}^{n}dx^{i}\otimes dx^{i}$ on $M=\mathbb{R}^{n}$. We denote its Levi-Civita connection by $\nabla^{g}$. The non-degenerate harmonic oscillator system [KKM2018] is given by the family of potentials
$V_{\mu}=\mu_{0}\sum_{i=1}^{n}(x^{i})^{2}+\sum_{i=1}^{n}\mu_{i}x^{i}+\mu_{n+1},$
where $\mu=(\mu_{0},\mu_{1},\dots,\mu_{n+1})\in\mathbb{R}^{n+2}$. It is an abundant non-degenerate irreducible conformally superintegrable system with the structure tensors given by, cf. [KSV2023],
[…]
It follows that $\bar{t}=0$ and $\hat{S}=0$, and hence the natural abundant Weylian structure is
$(M,\Phi,\hat{S})=(\mathbb{R}^{n},[(g,0)],0)\,,$
which corresponds to the Weyl structure $(\mathbb{R}^{n},[g],\nabla^{g})$. The associated semi-Weyl structure is thus given by
$(\mathbb{R}^{n},g,\nabla^{g},0)\,.$
Example 2.
Consider the Riemannian metric $g=\sum_{i=1}^{n}dx^{i}\otimes dx^{i}$ on $M=\mathbb{R}_{+}^{n}$, with Levi-Civita connection $\nabla^{g}$. The Smorodinski-Winternitz system is given by the family of potentials
$V_{\mu}=\mu_{0}\sum_{i=1}^{n}(x^{i})^{2}+\sum_{i=1}^{n}\frac{\mu_{i}}{(x^{i})^{2}}+\mu_{n+1},$
where $\mu=(\mu_{0},\mu_{1},\dots,\mu_{n+1})\in\mathbb{R}^{n+2}$. It is an abundant non-degenerate irreducible conformally superintegrable system with the structure tensors given by
[…]
where
$C=-\sum_{i=1}^{n}\frac{3}{x^{i}}dx^{i}\otimes dx^{i}\otimes\frac{\partial}{\partial x^{i}}\,.$
It follows that $t=-\frac{3}{n+2}\sum_{i=1}^{n}\ln(x^{i})$, up to an irrelevant constant [KSV2023]. Denoting the Kronecker delta by $\delta_{ij}$, we then define $S\in\Gamma(\mathrm{Sym}^{3}_{\circ}(T^{*}M))$ by
$S_{ijk}:=-\frac{3}{x^{i}}\delta_{ij}\delta_{jk}+\frac{3}{n+2}\frac{1}{x^{i}}\delta_{jk}+\frac{3}{n+2}\frac{1}{x^{j}}\delta_{ik}+\frac{3}{n+2}\frac{1}{x^{k}}\delta_{ij}$
and subsequently introduce $\hat{S}\in\Gamma(\mathrm{Sym}^{2}_{\circ}(T^{*}M)\otimes TM)$ by virtue of the inverse metric $g^{-1}$. The associated natural abundant Weylian structure of the Smorodinski-Winternitz system thus is
$\left(\mathbb{R}_{+}^{n},\left[\left(g,-\frac{3}{n+2}\sum_{i}\frac{dx^{i}}{x^{i}}\right)\right],\hat{S}\right)\,.$
It corresponds to the Weyl structure $(\mathbb{R}_{+}^{n},[g],D)$ where $D$ is the Levi-Civita connection of
$\exp\left(-\frac{2}{n+2}\sum_{j}\ln(x^{j})\right)\,g=\left(\prod_{j=1}^{n}x^{j}\right)^{-\frac{2}{n+2}}\,\sum_{i=1}^{n}dx^{i}\otimes dx^{i}.$
The underlying semi-Weyl structure is finally obtained as
$\left(\mathbb{R}_{+}^{n},g,D+\hat{S},\frac{2}{n+2}\sum_{i=1}^{n}\frac{dx^{i}}{x^{i}}\right)\,.$